April 26, 2011
All of the instruments in the room for STM were set at their former positions. LCD and CRT displays fell to the ground but these worked in the correct way.
STM-Raman(left), STM(center) and TCSPC(right) systems were restored.
May 13, 2011
Mr. Fujita and Mr. Sato succeeded in taking STM images of a HOPG surface (left) and a gold nano particle (right).
April 18, 2011
Mr. Shirasawa, Mr. Toyouchi and Mr. Kotaki restored nanosecond Nd:YAG lasers.
Mr. Kotaki checked his time-resolved Raman spectroscopy system (left). Second harmonic pulses irradiated a sample (right).
Femto Second Laser 1
April 1, 2011
A seed oscillator in a high power femtosecond laser system emitted light pulses.
April 14, 2011
A Nd:YLF pump laser radiated green light pulses (left) and an oscilloscope showed a regenerative amplifier waveform (center). Output power after a 2 path linear amplifier was 1.7 W at 2.5 kHz operation, and it's lower than it's before the earthquake. We should check and align all optical components.
Femto Second Laser 2
An oscilloscope showed another seed oscillator worked, but the oscillator was unstable.
Room for Optical Measurement
April 13, 2011
We cleaned the room for optical measurements.
Conventional UV-visible absorption and fluorescence spectroscopy systems were restored. Mr. Tomita tried taking absorption and fluorescence spectra of anthracene in ethanol.
Room for Chemicals
April 22, 2011
Broken glass ware was cleaned.
April 22, 2011
We cleaned our seminar room (left). Tap water and city gas supply started so we could boil water and make some coffee (right).
Each student has cleaned their desk and their room.